OpenStudy (anonymous):

A 12-ft ladder is leaning against a building. The bottom of the ladder is 7 ft from the building. How high is the top of the ladder?

OpenStudy (anonymous):

REMEMBER a^2+b^2=c^2 So b=7 and c=12 but we don't know 'a' yet. So the equation would be a^2+72=122 Find the square root of 7 and 12. So now the equation turns to a^2+49=144 Subtract 49 from 144: a^2+49=144

OpenStudy (anonymous):

* 144 - 49 - 49

OpenStudy (anonymous):

12*12=144 7*7=49 144-49=95\[\sqrt{95}=9.74\]So, x=9.74

OpenStudy (anonymous):

Well look you're left with a + 49 = 144 a = 144 - 49 a = 95 a = 9.75

OpenStudy (anonymous):

thanks guys!

OpenStudy (anonymous):

No problem. and good job @greenlegodude57

OpenStudy (anonymous):

@OrangeMaster \[\sqrt95 = 9.7467 = 9.75\]

OpenStudy (anonymous):

Thanks, you too.

OpenStudy (anonymous):

Same thing, just rounded differently. I'm sure he gets it anyhow.

OpenStudy (anonymous):

Yes.